Help me solve the system by substitution.
x - 10y = 2
x = 11y

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

if x=11y
the 11y-10y=2. That is substitution
y=2
x=22
(22, 2)

Answer by ikleyn(52788)

.
x - 10y =   2     (1)
x       = 11y     (2)


The second equation just expresses x = 11y  via y. 

So, substitute this expression into equation (2), replacing "x" there. You will get


11y - 10y = 2,    or

  y       = 2.


So, you just found  "y".  To find "x", substitute y= 2 into equation (2) to get

  x = 11*2 = 22.


ANSWER.  x= 22,  y= 2.


CHECK.  To check the solution and the answer, substitute the found values into the original equations.

        Do it on your own.

Solved and completed.